Grande03:

Jacqueline works at an aquarium where she keeps a record of the number of fish. The population of fish increases at a rate of 8.3% per year. The following expression represents the number of fish in the aquarium after t years. 21(1.083)^t The expression that reveals the approximate monthly growth rate of the population of fish can be modeled by the expression a^bt . Rounding to the nearest thousandth, determine the values of a and b that reveal the approximate monthly growth rate of the population of fish. A) a = 12 and b = 1.083 B) a = 1.006 and b = 12 C) a = 1.083 and b = 12 D) a = 12 and b = 1.006

crispyrat:

@AZ

Grande03:

?

crispyrat:

idk how to solve this but i think a user called AZ might so i @ him lol

Grande03:

oh ok thanks

AZ:

That's funny.. I answered this exact question earlier today

Grande03:

oh wow lol

AZ:

So we have an expression that is in the form of \( \Large a(1 + b)^t\) This tells us the growth or decline for every year (t is in years) to make it so that way we can calculate the growth/decay every month, we would have to do: \( \Large a\left(1 + \dfrac{b}{12}\right)^{12t}\)

AZ:

Well well well, I just checked and apparently you were the one who asked this question earlier LOL https://questioncove.com/study#/updates/60771e177bbef290ef155c00

Grande03:

I did?

AZ:

yes

Grande03:

It makes sense because my thing restarted but I don't remember this question

Grande03:

im sorry my memory is bad

AZ:

same numbers and all so we have \(\Large 21(1 + 0.083)^t\) that is what they originally gave us we want to make it into months so what is 0.083/12 = ?

AZ:

because basically all we're doing is \(\Large 21\left(1 + \dfrac{0.083}{12}\right)^{12t}\) and then 't' will be in months and it will be in the form of \(\Large a^{bt}\) so what would be the number in the parenthesis- that's going to be your 'a' and what is the number in the exponent right next to 't' that's your 'b'

Grande03:

thank uuuuu again lol

AZ:

you're welcome